(2x² + 15) + (29x - 45) = 180
1
Expert's answer
2012-10-22T07:28:36-0400
(2x² + 15) + (29x - 45) = 180
2x² + 15 + 29x - 45 = 180
2x² + 29x - 30 - 180=0
& 2x² + 29x - 210=0
D=29^2+4*210*2=2521>0 => there are two different real roots
x1,2=(-29+-sqrt(2521))/4
x1=(-29+sqrt(2521))/4
x2=(-29-sqrt(2521))/4